# IMAGHarmony: Controllable Image Editing with Consistent Object Quantity and Layout

### v1

@@ 4.1 Main Results
We quantitatively compare proposed IMAGHarmony with a broad range of state-of-the-art image editing methods across three cases: (1) mask-free editing approaches, including InstructPix2Pix [3] and MagicBrush [68]; (2) mask-based methods, such as LEDITS [57], LEDITS++ [2], and IC-Edit [69]; and (3) VLM-based approaches, including Show-o [61], SEED-X [11], and Emu3 [59].
@@
Figure 4: Qualitative comparisons with several state-of-the-art models on the HarmonyBench dataset.
@@
Quantitative Results. Table 1 summarizes the performance of IMAGHarmony compared to various state-of-the-art (SOTA) methods on the HarmonyBench dataset across three editing tasks: class, scene, and style editing. Our method consistently outperforms all SOTA methods across all metrics, including mask-free, mask-based, and VLM-guided approaches. Specifically, (1) compared to MagicBrush [68], the best-performing mask-free method, IMAGHarmony achieves improvements of 10.4% in OCA and 6.5% in AP on class editing, indicating stronger modeling of object quantity and spatial structure. Similar trends are observed in scene and style editing cases. (2) While the best mask-based method, IC-Edit [69], shows slight improvements over MagicBrush [68] in OCA and AP, benefiting from mask-based spatial priors, it still falls behind IMAGHarmony across all tasks. Notably, our framework achieves superior spatial control without any manual masks, thanks to the proposed HA attention module, which effectively captures quantity and layout representations. (3) Compared to VLM-guided methods, IMAGHarmony surpasses the strongest method, Show-o [61], by 13.8% and 14.3% in OCA and AP on scene editing. More importantly, it achieves significantly higher scores in IR, demonstrating superior semantic alignment. Unlike VLM-based approaches that typically require extensive parameters, IMAGHarmony delivers strong semantic understanding and controllability through its lightweight generative design. Overall, these results highlight the robustness and effectiveness of IMAGHarmony across diverse QL-Edit scenarios.
Qualitative Results. Figure 4 presents qualitative comparisons between IMAGHarmony and several SOTA methods under the QL-Edit setting. (1) In class editing (first three rows), IMAGHarmony accurately modifies object categories (e.g., “Three Apples” → “Three Pineapples”) while faithfully preserving both object count and spatial layout. In contrast, existing methods (e.g., MagicBrush [68], LEDITS [57]) often fail to change the category despite maintaining the correct number of objects. Furthermore, when editing instances with multiple quantities (e.g., “Seven Dogs” → “Seven Bears”), all methods produce incorrect counts and disrupt the original layout, whereas IMAGHarmony remains stable and structurally consistent. (2) In scene editing (rows four–five), our method successfully modifies the background (e.g., + “Fireworks in Background”) while preserving object integrity and spatial position. Other methods (e.g., LEDITS++ [2], IC-Edit [69]) often entangle foreground and background semantics, leading to failed scene modifications. Notably, SEED-X [11] additionally shows inconsistencies in object count. (3) In style editing (last two rows), IMAGHarmony transfers the target style (e.g., “Cyberpunk”) without altering object semantics or disrupting spatial arrangement. In contrast, existing methods frequently over-stylize or distort structure. For example, Show-o [61] mistakenly changes the object category during editing. In summary, IMAGHarmony effectively disentangles content from instructions, delivering semantically faithful and structurally consistent edits, even under complex multi-object.
@@
Figure 5: User study results on HarmonyBench. Our method achieves the best generation result.
@@
User Study. To assess the perceptual quality and semantic fidelity of the edited images, we conducted a user study involving 40 participants. We randomly sampled 50 prompts from the class, scene, and style editing tasks, along with corresponding outputs from all competing methods. Participants were asked to select the top-3 results for each prompt based on two criteria: semantic correctness (SC) [68] and overall quality (OQ) [3]. As shown in Figure 5, our method consistently achieved the highest scores across all metrics, indicating superior semantic alignment, perceptual realism, and user preference. These findings validate the effectiveness of IMAGHarmony from a human-centered evaluation perspective.
@@ 4.2 Ablation and Analysis
Effectiveness of the Module. To evaluate the effectiveness of each component in IMAGHarmony, we conduct ablation studies on four variants. Here, B0 removes both auxiliary textual features and our proposed HA module and PNS strategy, injecting source image features via whole IP attention [67]; B1 also excludes auxiliary text and PNS, but injects the source image only through a layout-specific attention layer, similar to InstantStyle [58]; B2 retains PNS but removes the HA module, thus lacking explicit modeling of object quantity and implicit layout reasoning; B3 removes the PNS strategy while keeping HA, using a random seed during inference.
Table 2: Ablation study on HarmonyBench.
Methods OCA ($\uparrow$) AP ($\uparrow$) IR ($\uparrow$)
B0 43.2 21.7 0.161
B1 48.6 56.9 0.193
B2 51.3 49.9 0.367
B3 71.3 68.8 0.322
@@
As shown in Table 2, B0 performs the worst across all metrics, highlighting the need for structure-aware and semantically guided modeling. B1 improves AP (21.7% → 56.9%) via layout-specific injection but still lags in OCA and IR, suggesting limited gains from shallow spatial cues. B2 introduces PNS, yielding better OCA and IR, yet lacks explicit modeling of object quantity and layout. B3 benefits from the HA module, achieving higher OCA and AP , but suffers in IR due to unstable seed initialization. In contrast, the full IMAGHarmony model with both HA and PNS, achieves the best performance, confirming the complementary benefits of harmony features and seed-aware inference.
Figure 6: Attention map comparison. (a) w/o HA. (b) w/ HA. HA module better captures object count and layout.
Visualization of the Feature Map. To further validate the effectiveness of the proposed HA module in modeling object quantity and spatial layout, we visualize and compare attention distributions with and without the HA module. From Figure 6, the heatmaps are generated by averaging activation maps from the "Down4" blocks of the UNet. The source image contains multiple well-distributed objects, representing a typical multi-object editing scenario. In Figure 6 (a), without the HA module (w/o HA) exhibits dispersed attention with weak structural focus, often leading to missing or misaligned objects. In contrast, Figure 6 (b) shows that our model with HA produces sharper, more concentrated attention aligned with the object regions. These visualizations demonstrate that the HA module effectively enhances the model’s ability to capture object count and spatial structure, thereby improving semantic and structural consistency in QL-Edit.
Injection Location Analysis. As shown in Figure 7, we evaluate the impact of injecting the harmony feature at different locations within the UNet, including downsampling (D), middle (M), and upsampling (U) blocks. D1 refers to the first downsampling block. The results demonstrate that inserting the harmony feature at the fourth downsampling block (D4), yields the highest performance across both object count accuracy (OCA) and average precision (AP), significantly outperforming all other configurations. This suggests that introducing structure-aware guidance at D4 provides a better balance between semantic and spatial structure, allowing the model to capture quantity and layout consistency more effectively. In contrast, injecting at shallower or deeper layers fails to model structural relationships, resulting in inferior performance.
@@
Impact of Hyperparameters. We further study the impact of two key hyperparameters, $\alpha$ and $\beta$, which control the influence of the HA module and IP Attention modules, respectively, as defined in Eq. (4) and Eq. (5). We vary each parameter independently from 0.0 to 1.0 in increments of 0.2, while keeping the other fixed at 1.0. As shown in Figure 8, the best editing quality occurs when both $\alpha=1.0$ and $\beta=1.0$. In contrast, lower values lead to unstable object quantity and spatial layout, resulting in reduced structural fidelity and semantic consistency.

### v2

@@ 4.2 Main Results
We quantitatively compare proposed IMAGHarmony with a broad range of state-of-the-art image editing methods across three cases: (1) mask-free editing approaches, including InstructPix2Pix [19] and MagicBrush [18]; (2) mask-based methods, such as LEDITS [23], LEDITS++ [25], LoMOE [68], and IC-Edit [26]; and (3) VLM-based approaches, including Show-o [59], SEED-X [58], and Emu3 [56].
@@
Figure 4: Qualitative comparisons with several state-of-the-art models on the HarmonyBench dataset. A red cross indicates an unsuccessful edit, where the output shows no difference compared to the source image.
@@
Quantitative Results. TABLE III summarizes the performance of IMAGHarmony compared to various state-of-the-art (SOTA) methods on the HarmonyBench dataset across three editing tasks: class, scene, and style editing. Our method consistently outperforms all SOTA methods across all metrics, including mask-free [19, 18], mask-based [23, 25, 68, 26], and VLM-guided approaches [59, 58, 56]. Specifically, (1) compared to MagicBrush [18], the best-performing mask-free method, IMAGHarmony achieves improvements of 10.4 in OCA and 6.5 in AP on class editing, indicating stronger modeling of object quantity and spatial structure. Similar trends are observed in scene and style editing cases. (2) While the best mask-based method, IC-Edit [26], shows slight improvements over MagicBrush [18] in AP and IR on average, benefiting from mask-based spatial priors, it still falls behind IMAGHarmony across all tasks. Notably, our framework achieves superior spatial control without any manual masks, thanks to the proposed HA module, which effectively captures quantity and layout representations. (3) Compared to VLM-guided methods, IMAGHarmony surpasses the strongest method, Show-o [59], by 13.8 and 14.3 in OCA and AP on scene editing. More importantly, it achieves significantly higher scores in IR, demonstrating superior semantic alignment. Unlike VLM-based approaches that typically require extensive parameters, IMAGHarmony delivers strong semantic understanding and controllability through its lightweight generative design.
Furthermore, we evaluate HarmonyBench by stratifying images according to the number of objects to be edited, and report the results in TABLE IV. Most SOTA methods degrade substantially as object cardinality increases: for example, MagicBrush [18] drops from 89.1/87.8 (OCA/AP) at 1–2 objects to 78.3/71.2 at more than 9 objects, and IC-Edit [26] falls from 89.3/88.1/0.402 to 78.8/74.3/0.294; LEDITS++ [25] even yields negative IR in dense scenes, indicating a breakdown of structural/layout fidelity. In contrast, the proposed IMAGHarmony is outstandingly stable when the number of objects changes: the SD1.5 variant remains nearly flat (e.g., IR 0.414→0.408 from 1–2 to 9+), and the SDXL variant is stronger overall with only minor reductions (e.g., OCA 89.9→89.2, AP 88.5→85.1). Moreover, as shown in TABLE III and TABLE IV, IMAGHarmony achieves highly competitive results on both SD1.5 and SDXL backbones, demonstrating strong generalization and plug-and-play capability. Overall, these results highlight the robustness and effectiveness of IMAGHarmony across diverse QL-Edit scenarios.
@@
Figure 5: User study results on HarmonyBench. Our method receives the highest praise from users.
Qualitative Results. Fig. 4 presents qualitative comparisons between IMAGHarmony and several SOTA methods under the QL-Edit setting. (1) In class editing (first three rows), IMAGHarmony accurately modifies object categories (e.g., "Three Apples" → "Three Pineapples") while faithfully preserving both object count and spatial layout. In contrast, existing methods (e.g., MagicBrush [18], LEDITS [23]) often fail to change the category despite maintaining the correct number of objects. Furthermore, when editing instances with multiple quantities (e.g., "Seven Bears" → "Seven Lions"), all methods produce incorrect counts and disrupt the original layout, whereas IMAGHarmony remains stable and structurally consistent. (2) In scene editing (rows four–five), our method successfully modifies the background (e.g., + "in Fireworks") while preserving object integrity and spatial position. Other methods (e.g., LEDITS++ [25], IC-Edit [26]) often entangle foreground and background semantics, leading to failed scene modifications. Notably, SEED-X [58] additionally shows inconsistencies in object count. (3) In style editing (last two rows), IMAGHarmony transfers the target style (e.g., +"Cyberpunk City Style") without altering object semantics or disrupting spatial arrangement. In contrast, existing methods frequently over-stylize or distort structure. For example, Show-o [59] mistakenly changes the object category during editing. In summary, IMAGHarmony effectively disentangles content from instructions, delivering semantically faithful and structurally consistent edits, even under complex multi-object.
@@
User Study. We conducted a user study with 40 participants to assess IMAGHarmony from a human-centered perspective. Fifty prompts were randomly sampled across the three QL-Edit tasks (class, scene, style), and edited results from all competing methods (including IMAGHarmony) were collected. For each prompt, participants selected the top-3 outputs under two criteria: semantic correctness (SC; object category, count, and layout) [18] and overall quality (OQ; coherence and aesthetics) [19]. To reduce bias, results were presented in randomized, anonymized order, and each participant evaluated a distinct subset of prompt–method pairs for balanced coverage. As shown in Fig. 5, IMAGHarmony consistently achieved the highest scores on both SC and OQ, indicating superior semantic alignment, perceptual realism, and user preference. These findings validate the effectiveness of IMAGHarmony from a human-centered evaluation standpoint.
TABLE V: Ablation study of the HA module and PNS strategy on HarmonyBench.
[…]
Baseline 43.2 21.7 0.161
B1 48.6 56.9 0.193
B2 84.6 82.4 0.366
B3 85.1 82.7 0.369
B4 86.7 84.4 0.391
IMAGHarmony 88.8 85.9 0.408
@@ 4.3 Ablation and Analysis
Effectiveness of the Module. To evaluate the contribution of each component (i.e., the HA module and the PNS strategy) in IMAGHarmony, we conduct ablations over five variants (Baseline and B1–B4):
  • Baseline: no HA module and no PNS strategy; source-image features are injected via whole cross attention.
  • B1: use only one cross attention layer within the HA module; no PNS strategy.
  • B2: use the full HA module; no PNS strategy.
  • B3: use the full HA module; apply PNS strategy only at inference.
  • B4: use the full HA module; apply PNS strategy only during training.
As shown in TABLE 4.2, the Baseline is lowest on all three metrics (OCA 43.2 / AP 21.7 / IR 0.161). Using only a single cross attention within HA module (B1) markedly increases AP (+35.2 to 56.9) but yields limited improvements in OCA and IR (+5.4 / +0.032), indicating that cross-modal interaction at this layer can enhance spatial alignment to some extent but is insufficient for stable modeling of quantity–layout relations. When the full HA module (with QL attention and cross attention, B2) is enabled, all three metrics rise substantially (OCA 84.6, AP 82.4, IR 0.366), showing that explicit quantity modeling together with implicit layout reasoning is the primary source of gains. Applying PNS strategy only at inference (B3) provides little additional benefit over B2 (OCA +0.5, AP +0.3, IR +0.003), whereas introducing PNS strategy during training (B4) leads to further improvements over B3 (OCA +1.6, AP +1.7, IR +0.022), indicating that seed diversity exposure during training is more critical for trajectory stability and structural fidelity. Finally, the full model, using PNS strategy in both training and inference together with HA, achieves the best results (OCA 88.8 / AP 85.9 / IR 0.408). Notably, AP continues to increase from B2 through B4 to Ours, consistent with the qualitative visualizations, suggesting that the combination of HA module and PNS strategy not only improves object counting accuracy but also significantly strengthens layout consistency and overall image quality.
Figure 6: Attention map comparison. (a) w/o HA module. (b) w/ HA module. HA module better captures object count and layout.
Visualization of the Feature Map. To further validate the effectiveness of the proposed HA module in modeling object quantity and spatial layout, we visualize and compare attention distributions with and without the HA module. From Fig. 6, the heatmaps are generated by averaging activation maps from the "Down4" blocks of the UNet. The source image contains multiple well-distributed objects, representing a typical multi-object editing scenario. In Fig. 6 (a), without the HA module (w/o HA module) exhibits dispersed attention with weak structural focus, often leading to missing or misaligned objects. In contrast, Fig. 6 (b) shows that our model with HA module produces sharper, more concentrated attention aligned with the object regions. These visualizations demonstrate that the HA module effectively enhances the model’s ability to capture object count and spatial structure, thereby improving semantic and structural consistency in QL-Edit.
Injection Location Analysis. As shown in Fig. 7, we evaluate the impact of injecting the harmony feature at different locations within the UNet, including downsampling (D), middle (M), and upsampling (U) blocks. D1 refers to the first downsampling block. The results demonstrate that inserting the harmony feature at the fourth downsampling block (D4) yields the highest performance across both object count accuracy (OCA) and average precision (AP), significantly outperforming all other configurations. This suggests that introducing structure-aware guidance at D4 provides a better balance between semantic and spatial structure, allowing the model to capture quantity and layout consistency more effectively. In contrast, injecting at shallower or deeper layers fails to model structural relationships, resulting in inferior performance. These findings validate the effectiveness and necessity of injecting harmony features at the D4 stage for structurally consistent image editing. Moreover, the D4 setting introduces minimal computational overhead in practice, keeping training and inference latency nearly unchanged. These findings validate the effectiveness and necessity of injecting harmony features at the D4 stage for structurally consistent image editing.
Feature Fusion Strategies. The preceding experiments show that the quantity and layout (QL) attention explicitly injects quantity–layout priors, strengthens structural consistency and semantic alignment, and reduces duplication/omission and off-layout drift in multi-object scenes, yielding more stable OCA/AP and IR. Building on this, we further compare and analyze QL attention against three representative feature-fusion strategies, i.e., Q-Former [76], MLP fusion [77], and gated attention [78], and conduct both quantitative and qualitative evaluations on the HarmonyBench dataset. Specifically, Q-Former leverages learnable queries to extract salient information from auxiliary text and interact with visual features; MLP fusion concatenates multimodal embeddings and passes them through multilayer perceptrons for integration; and Gated Attention employs gating mechanisms to modulate the contributions of different modalities. In contrast, our HA module explicitly models object quantity and implicitly captures spatial layout, providing stronger structural awareness and cross-modal alignment.
Figure 8: Qualitative results of different fusion strategies on the HarmonyBench dataset.
TABLE VI: Quantitative results of different fusion strategies on the HarmonyBench dataset.
[…]
Q-Former [76] 42.3 41.9 0.134
MLP [77] 44.6 36.8 0.078
Gated Attention [78] 67.4 55.9 0.194
@@
[…]
@@
Impact of Hyperparameters. We further study the impact of two key hyperparameters, $\alpha$ and $\beta$, which control the influence of the HA module and IP Attention modules, respectively, as defined in Eq. (4) and Eq. (5). We vary each parameter independently from 0.0 to 1.0 in increments of 0.2, while keeping the other fixed at 1.0. As shown in Fig. 9, the best editing quality occurs when both $\alpha=1.0$ and $\beta=1.0$. Furthermore, we observe that increasing $\alpha$ progressively heightens IMAGHarmony’s sensitivity to the quantity attribute, whereas increasing $\beta$ progressively heightens its sensitivity to the category attribute. This trend is consistent with our design: the HA module controlled by $\alpha$ injects quantity–layout priors and therefore primarily affects counting and structural consistency; the IP Attention controlled by $\beta$ emphasizes semantic alignment and thus is more sensitive to category discrimination. In contrast, lower values lead to unstable object quantity and spatial layout, resulting in reduced structural fidelity and semantic consistency.
TABLE VII: Quantitative comparisons on MagicBrush dataset. IMAGHarmony achieves the best performance.
[…]
SEED-X [58] 40.8 21.4 0.337
Show-o [59] 51.2 51.9 0.114
Emu-3 [56] 58.2 47.0 0.187
LEDITS [23] 67.2 63.9 0.017
LEDITS++ [25] 74.8 72.5 0.093
LoMOE [68] 82.3 78.3 0.105
InstructPix2Pix [19] 85.4 78.7 0.314
MagicBrush [18] 88.2 77.1 0.345
IC-Edit [26] 88.6 84.8 0.436
Ours 91.7 89.1 0.479
Figure 10: Qualitative results on editing two object categories simultaneously compared with SOTA methods.
4.4 More Results and Analysis
Results of MagicBrush Dataset. To further assess generalization, we evaluate on the widely used MagicBrush [18] dataset. As shown in TABLE VII, our method achieves the best scores on all metrics (OCA 91.7, AP 89.1, IR 0.479). Compared with the strongest SOTA method, IC-Edit [26] (88.6, 84.8, 0.436), the gains are +3.1, +4.3, and +0.043, indicating simultaneous improvements in object-count accuracy, spatial alignment, and structural stability. Relative to MagicBrush [18] itself (88.2, 77.1, 0.345), the increases in AP and IR are particularly notable (+12.0 and +0.134), suggesting better preservation of the reference layout while performing the edits. Other methods exhibit one-dimensional strengths with overall limitations, for example, LoMOE [68] attains a relatively high OCA (82.3) but a low IR (0.105), and LEDITS++ [25] reaches only 72.5 AP, limiting global structural consistency. Overall, these results substantiate that the quantity–layout priors injected by HA module and the seed robustness provided by PNS strategy generalize to a mainstream benchmark and offer complementary, additive gains.
Figure 11: Qualitative results with a fixed source image. IMAGHarmony produces more stable and consistent images.
TABLE VIII: Analysis of more variants. Higher is better for all metrics.
[…]
(1) VLM Evaluator
VLM GPT-4V [79] 88.7 85.8 0.407
CogVLM2 [80] 88.8 85.9 0.408
[.6pt/2pt]               (2) Effect of Candidate Seeds
K 3 87.4 83.9 0.386
5 88.8 85.9 0.408
10 88.8 86.0 0.411
[.6pt/2pt]               (3) Shallow Denoising Steps
Steps 10 88.8 85.9 0.408
20 88.8 85.9 0.412
30 89.8 86.2 0.414
Results on Multi-Class Editing. Fig. 10 presents qualitative results comparing IMAGHarmony with SOTA methods on editing two object categories simultaneously, an especially challenging scenario due to the need for fine-grained control over both object count and category consistency. For instance, the prompt "Two Dogs and Two Cats" is edited to "Two Tigers and Two Lions," and "Three Apples and Two Eggs" is edited to "Three Oranges and Two Vases." The VLM-based approaches (e.g., SEED-X [58] and Show-o [59]) achieve better semantic grounding through large language models but still suffers from object blending and inconsistent quantities, especially when the instruction involves more than one distinct category. Then, MagicBrush [18] tends to overfit to one dominant category, often failing to modify both targets accurately. It also struggles to preserve the spatial arrangement and object count, frequently producing duplicated instances. Lastly, IC-Edit [26] offers slightly improved layout retention by leveraging structural priors, but still exhibits semantic confusion, with mixed or partially transformed objects. Moreover, it causes serious interference with the background. In contrast, IMAGHarmony demonstrates superior performance by accurately transforming all specified object categories while strictly preserving their respective counts and spatial layout. This is attributed to its harmony aware representation and preference-guided noise selection, which together enable fine-grained structural and semantic alignment even in complex multi-category editing tasks. These results highlight the robustness and generalization ability of IMAGHarmony in multi-object scenarios beyond single-category edits.
Results with a Fixed Source Image. To further assess generalization, we perform both class editing and style editing on the same source image. As illustrated in Fig. 11, class editing converts "Sheep" into "Dogs" and "Crows", and converts "Lions" into "Wolves" and "Pigs"; style editing renders the former in +"Cyberpunk" and +"Watercolor" styles, and the latter in +"Futuristic" and "Inkpainting" styles. For class editing, our method faithfully follows each semantic instruction while preserving the original object counts and spatial layout; the generated images reflect the target categories without altering the underlying structure of the reference, yielding realistic and coherent edits. In contrast, SOTA methods [58, 59, 18, 26] often exhibit noticeable errors under different instructions, including incorrect category replacement and mismatched object counts. For style editing, IMAGHarmony successfully conveys the target styles while maintaining the original object quantity, categories, and spatial layout; the visual outputs show strong style fidelity with minimal distortion, and the scene structure remains intact across all variations. By comparison, existing methods [58, 59, 18, 26] tend to suffer from over-stylization or structural collapse under strong aesthetic transformations. For example, IC-Edit [26] often fails under +"Futuristic and +"Inkpainting" prompts, failing to follow the editing instructions. These observations highlight the advantage of IMAGHarmony in decoupling structural modeling from style modulation, enabling semantically faithful and layout-consistent edits across diverse style prompts.
TABLE IX: Comparison of training datasets, trainable parameters, generation speed, and model scale of the proposed method with recent state-of-the-art methods.
[…]
Training Data ($\downarrow$) 158.0M 2.0B 8.8k 50.0K 0.2K
Train. Params ($\downarrow$) 2.6B 1.3B 1.1B 214.0M 10.6M
Speed ($\downarrow$) 12.4s 10.4s 24.2s 9.7s 6.6s
Model Scale ($\downarrow$) 13B 1.3B 1.1B 12B 0.78B
Analysis of More Variants. TABLE 4.4 reports the effects of three factors. First, using GPT-4V [79] versus CogVLM2 [80] as the seed-scoring evaluator yields nearly identical results (e.g., OCA 88.7 vs. 88.8, AP 85.8 vs. 85.9, IR 0.407 vs. 0.408), indicating that the PNS strategy is robust to the choice of VLM and generalizes well. Second, enlarging the candidate pool $K$ brings modest gains (e.g., IR: 0.386 $\rightarrow$ 0.408 $\rightarrow$ 0.411) but saturates beyond $K=5$ while inference cost grows roughly linearly with $K$; we therefore choose $K=5$ as a practical trade-off. Finally, increasing the shallow denoising steps provides marginal improvements (IR: 0.408 $\rightarrow$ 0.412 $\rightarrow$ 0.414; with slight increases in OCA/AP) at the expense of additional runtime; we adopt 10 steps as a balanced configuration. In summary, we adopt CogVLM2 as the VLM evaluator, set the candidate pool to $K=5$, and use $10$ denoising steps.
[…]